# Interpreting Key Mechanisms of Factual Recall in Transformer-Based Language Models

### v1

@@ 4.1 Comparing Zero-shot with Few-shot Scenarios
We construct two new datasets: one with a single in-context demonstration (one-shot) and another with two demonstrations (two-shot). We achieve this by appending one or two prefixes in the form of “The capital of $X_{\text{ICL}}$ is $Y_{\text{ICL}}$,” respectively, to the original prompts. the pair ($X_{\text{ICL}}$, $Y_{\text{ICL}}$) is randomly selected and distinct from ($X$, $Y$).
@@
Using these datasets, we conduct “patching $\textbf{a}^{l,h}_{t=-1}\rightarrow\textbf{r}^{11}_{\text{post},t=-1}$” experiments as described in §3.2. We also plot the probability dynamics of $X$ and $Y$ across the layers. The results are shown in Figure 5. Upon comparing the results of 0-shot and few-shot experiments, we observe that the mechanisms studied in the zero-shot scenario persist in the few-shot scenarios. This persistence manifests in two aspects: Firstly, the consistency of probability dynamics and influential heads across varying numbers of ICL demonstrations; Secondly, the consistent decline in the probability of both $X$ and $Y$ as they reach the final layers across all scenarios. Now, it becomes evident that the identification of the “saturation” phase by Merullo et al. (2023) is not due to in-context demonstrations but rather stems from the metric they utilize, the token rank rather than detailed probability. Consequently, the final declines are overlooked. The following subsections will delve into the reasons behind this universal decline at the final layer.
ICL Improves Model Confidence in Generating Capitals’ Name
@@
Before delving into the universal decline, there is an extra crucial finding when comparing the zero-shot results to the few-shot results. One noticeable distinction lies in the emergence of probabilities for $X$ and $Y$ at layer 8 in few-shot settings, compared to their emergence until layer 9 in 0-shot settings. This distinction can be attributed to L8H11 and L8H6. As the number of ICL demonstrations increases, these heads become more influential. L8H11 is identified as a “mover” head, while L8H6 serves as an “inhibition” head, as described by (Wang et al., 2023). The inhibition head impacts subsequent mover heads’ attention pattern. We have discussed how to identify a mover head, and for detecting an inhibition head, please refer to Wang et al. (2023).
@@
The ICL demonstrations enhance the model’s confidence in generating the name of capitals from two key aspects: Firstly, an additional mover head, L8H11, moves the country names to the final position, implying that with more ICL demonstrations, the model gains insights into the “argument” earlier, allowing for deeper processing of the answer across more layers. Secondly, in the zero-shot scenario, the model tends to produce “safe” tokens such as “not” to follow the prompt, resulting in sentences like “The capital of $X$ is not…” which, although coherent, lack useful information. This might stem from a lack of confidence in the output domain or format. However, in few-shot settings, L8H6 pays considerable attention from the final token to $Y_{\text{ICL}}$, i.e., the in-context capital names. It may prompt the model to recognize that the expected answer is the name of a capital city.
4.2 Negative heads Suppress $X$
@@ 4.3 The Final Layer Avoids Overconfidence
Despite our efforts to block out each attention head, the suppression effect persists, suggesting that a large portion of all heads may contribute together to the suppression of $Y$. Visualizing the attention patterns of each head (see Figure 10) provides more insights. It becomes evident from the figure that most heads “park” their attention on the initial token. This pattern is consistent across various tasks and models, indicating its generality. Previous works (Miller, 2023; Xiao et al., 2023) suggest that this attention pattern indicates that attention heads are inactive and indifferent to the context. We analyzed the output of the OV circuit in these heads at the initial position and observed that some of them output frequent English tokens such as "the," ",", ".", "a," and "and," while others produce letters and some subwords irrelevant to the task. We hypothesize that these inactive heads move information in the initial position to the final position, diluting the content of $Y$ in the residual stream. To test our hypothesis, we applied attention masks to restrict these heads to attend solely to the final position (i.e., only the diagonal is unmasked). The result, as depicted in Figure 12(a.2), (b.2), and (c.2), shows a mitigation of suppression at $\textbf{r}^{11}_{\text{mid}}$. For example, with zero-shot, We improved the probability of $Y$ at $\textbf{r}^{11}_{\text{mid}}$ from 13.41% to 31.68%.
@@
Further, we delve into examining the learned intercept of the final layer to gain deeper insights into the workings of the final MLP. We hypothesize that the final MLP aligns the residual stream with $\mathbb{E}[\textbf{W}_{U}]$, the expectation of output logits over the training corpus, thereby effectively mitigating significant training losses. $\mathbb{E}[\textbf{W}_{U}]$ is computed as the average of $w_{v}\cdot\textbf{W}_{U}[v]$ over $V$ tokens. Here, $w_{v}$ represents the relative word frequency derived from OpenWebText (Gokaslan and Cohen, 2019), the training corpus of GPT-2. We utilize the word frequency data provided by Kobayashi et al. (2023).
@@
We acknowledge that the hypothesis regarding the final MLP layer remains anecdotal since we have only tested it within the GPT-2 family due to a lack of training corpus data for other models. Nevertheless, subtracting the intercept from the residual stream in the final layer empirically proves to be an effective approach in mitigating the suppression of correct predictions. Results on GPT-2 small are depicted in Figure 12(a.3), (b.3), and (c.3). Through this approach, for example, we were able to enhance the zero-shot probability of $Y$ at $\textbf{r}^{11}_{\text{post}}$ in GPT-2 small from $15.51\%$ to $25.93\%$. The efficacy of two anti-suppression techniques—namely, applying attention masks and subtracting intercepts—has been validated across various models and tasks. Please refer to the appendix D and E for further details.
5 Related Works
@@
In investigating the factual recall of language models, Geva et al. (2023, 2022) traced the information change in factual recall tasks. Yu et al. (2023); Ortu et al. (2024) studied important attention heads in a counterfactual setting. (Heinzerling and Inui, 2024) is a contemporary work that finds the specific direction in activation space encodes the semantics of model final outputs. None of these works provide an in-depth analysis of the mechanisms behind each module’s function. Merullo et al. (2023) described the factual recall as a process of “passing an argument and then applying the function.” Based on their analogy, we take a further step to provide the details of how “arguments” are identified from the context and passed to the “function.” Similar to our paper, Geva et al. (2023) also found that the OV matrix of attention heads contains knowledge. Other related works focus on factual recall and mechanistic interpretability. Our findings do not overlap with theirs but complement each other. Lieberum et al. (2023) conducted a multiple-choice task involving factual knowledge within large language models. Yu and Ananiadou (2024) pinpointed the location of knowledge within neurons in MLPs and interpreted how the residual stream activates knowledge in MLPs.
@@
Elhage et al. (2021) proposed that Transformers operate using a “memory management” approach, where modules are reading from and writing to the residual stream. Dao et al. (2023) trained a 4-layer model and found that certain attention heads could erase the outputs of previous attention heads. McGrath et al. (2023) observed the erasure of tokens by MLPs. More than observations, we dissected the role of MLPs by quantifying their impact on individual heads, and interpreting their role in “function application.”
6 Conclusion
In this paper, we conduct a thorough investigation into the factual recall mechanisms utilized by language models. We studied the mechanisms behind “argument passing” and “function application”. Furthermore, we uncovered the prevalence of anti-overconfidence mechanisms in language models. Our proposed analysis method, based on linear regression, effectively decomposes MLP outputs into components that are easily understandable to humans. This method has been substantiated through numerous empirical experiments and lays a valuable foundation for our interpretations. Finally, we study the anti-overconfidence mechanisms in the final layer, which is a universal mechanism regardless of models, tasks, and the number of in-context demonstrations.
Acknowledgments and Disclosure of Funding
Appendix A Details of Datasets
@@
In addition to the country-capital task, we also conduct experiments on a product-developer task. Table 4 shows prompt templates we used. Here are Product-Developer ($X$, $Y$) pairs we used: (iPhone, Apple), (Windows7, Microsoft), (GTX1060, Nvidia), (YouTube, Google), (Firefox, Mozilla), (VirtualBox, Oracle), (Instagram, Facebook), (Pentium, Intel), (Steam, Valve), (Radeon, AMD), (Photoshop, Adobe), (Boeing747, Boeing), (PlayStation, Sony), (Kindle, Amazon), (GameBoy, Nintendo).
@@ 7 Keep in mind, the capital of $X$ is
8 Just a reminder, the capital of $X$
9 As we all know, the capital of $X$
10 According to the textbook, the capital of $X$
@@ 14 To emphasize, $X$ is developed by
Appendix B Details on Solving the Linear Regression

### v4

@@ A.1 Zero-shot mechanisms still work in few-shot scenarios
We construct two new datasets: one with a single in-context demonstration (one-shot) and another with two (two-shot). We achieve this by appending one or two prefixes in the form of “The capital of $X_{\text{ICL}}$ is $Y_{\text{ICL}}$,” respectively, to the original prompts. each ($X_{\text{ICL}}$, $Y_{\text{ICL}}$) is randomly selected and distinct from ($X$, $Y$).
@@
Using these datasets, we conduct “patching $\textbf{a}^{l,h}_{t=-1}\rightarrow\textbf{r}^{11}_{\text{post},t=-1}$” experiments as described in §3.2. We also plot the probability dynamics of $X$ and $Y$ across the layers. The results are shown in Figure 5. Upon comparing the results of 0-shot and few-shot experiments, we observe that the mechanisms studied in the zero-shot scenario persist in the few-shot scenarios. This persistence manifests in two aspects: Firstly, the consistency of probability dynamics and influential heads across varying numbers of ICL demonstrations; Secondly, the consistent decline in the probability of both $X$ and $Y$ as they reach the final layers across all scenarios. Now, it becomes evident that the identification of the “saturation” phase by Merullo et al. [21] is not due to in-context demonstrations but rather stems from the metric they utilize, the token rank rather than detailed probability. Consequently, the final declines are overlooked. The following subsections will delve into the reasons behind this universal decline at the final layer.
A.2 ICL improves model confidence in generating capitals’ name
@@
Before delving into the universal decline, there is an extra crucial finding when comparing the zero-shot results to the few-shot results. One noticeable distinction lies in the emergence of probabilities for $X$ and $Y$ at layer 8 in few-shot settings, compared to their emergence until layer 9 in 0-shot settings. This distinction can be attributed to L8H11 and L8H6, where L8H11 is identified as a “mover” head. As the number of ICL demonstrations increases, these heads become more influential.
@@
The ICL demonstrations enhance the model’s confidence in generating the name of capitals from two key aspects: Firstly, an additional mover head, L8H11, moves the country names to the final position, implying that with more ICL demonstrations, the model gains insights into the “argument” earlier, allowing for deeper processing of the answer across more layers. Secondly, in the zero-shot scenario, the model tends to produce “safe” tokens such as “not” to follow the prompt, resulting in sentences like “The capital of $X$ is not…” which, although coherent, lack useful information. This might stem from a lack of confidence in the output domain or format. However, in few-shot settings, L8H6 pays considerable attention from the final token to $Y_{\text{ICL}}$, i.e., the in-context capital names. It seems prompt the model to recognize that the expected answer is the name of a capital city. We leave the exploration of the following questions to future research, as the answer is non-trivial: How do ICL demonstrations trigger the earlier activation of mover heads? How does the model determine that it should output city names?
A.3 Negative heads suppress $X$
A.4 The final layer avoids overconfidence
@@
Despite our efforts to block out each attention head, the suppression effect persists, suggesting that a large portion of all heads may contribute together to the suppression of $Y$. Visualizing the attention patterns of each head (see Figure 8) provides more insights. It becomes evident from the figure that most heads “park” their attention on the initial token. This pattern is consistent across various tasks and models, indicating its generality. Previous works [23, 42] suggest that this attention pattern indicates that attention heads are inactive and indifferent to the context. We analyzed the output of the OV circuit in these heads at the initial position and observed that some of them output frequent English tokens such as "the," ",", ".", "a," and "and," while others produce letters and some subwords irrelevant to the task. We hypothesize that these inactive heads move information in the initial position to the final position, diluting the content of $Y$ in the residual stream. To test our hypothesis, we applied attention masks to restrict these heads to attend solely to the final position (i.e., only the diagonal is unmasked). The result, as depicted in Figure 9(a.2), (b.2), and (c.2), shows a mitigation of suppression at $\textbf{r}^{11}_{\text{mid}}$. For example, with zero-shot, We improved the probability of $Y$ at $\textbf{r}^{11}_{\text{mid}}$ from 13.41% to 31.68%.
@@
Further, we delve into examining the learned intercept of the final layer to gain deeper insights into the workings of the final MLP. We hypothesize that the final MLP aligns the residual stream with $\mathbb{E}[\textbf{W}_{U}]$, the expectation of output logits over the training corpus, thereby effectively mitigating significant training losses. $\mathbb{E}[\textbf{W}_{U}]$ is computed as the average of $w_{v}\cdot\textbf{W}_{U}[v]$ over $V$ tokens. Here, $w_{v}$ represents the relative word frequency derived from OpenWebText [11], the training corpus of GPT-2. We utilize the word frequency data provided by [14].
@@
We acknowledge that the hypothesis regarding the final MLP layer remains anecdotal since we have only tested it within the GPT-2 family due to a lack of training corpus data for other models. Nevertheless, subtracting the intercept from the residual stream in the final layer empirically proves to be an effective approach in mitigating the suppression of correct predictions. Results on GPT-2 small are depicted in Figure 9(a.3), (b.3), and (c.3). Through this approach, for example, we were able to enhance the zero-shot probability of $Y$ at $\textbf{r}^{11}_{\text{post}}$ in GPT-2 small from $15.51\%$ to $25.93\%$. The efficacy of two anti-suppression techniques—namely, applying attention masks and subtracting intercepts—has been validated across various models and tasks. Please refer to Appendix F, G and H for further details.
Appendix B Related works
@@
In investigating the factual recall of language models, Geva et al. [9, 10] traced the information change in factual recall tasks. Yu et al. [43] and Ortu et al. [28] studied important attention heads in a counterfactual setting. [12] is a contemporary work that finds the specific direction in activation space encodes the semantics of model final outputs. None of these works provide an in-depth analysis of the mechanisms behind each module’s function. Merullo et al. [21] described the factual recall as a process of “passing an argument and then applying the function.” Based on their analogy, we take a further step to provide the details of how “arguments” are identified from the context and passed to the “function.” Similar to our paper, Geva et al. [9] also found that the OV matrix of attention heads contains knowledge. Other related works focus on factual recall and mechanistic interpretability. Our findings do not overlap with theirs but complement each other. Lieberum et al. [17] conducted a multiple-choice task involving factual knowledge within large language models. Yu et al. [44] pinpointed the location of knowledge within neurons in MLPs and interpreted how the residual stream activates knowledge in MLPs.
@@
Elhage et al. [8] proposed that Transformers operate using a “memory management” approach, where modules are reading from and writing to the residual stream. Dao et al. [4] trained a 4-layer model and found that certain attention heads could erase the outputs of previous attention heads. McGrath et al. [19] observed the erasure of tokens by MLPs. More than observations, we dissected the role of MLPs by quantifying their impact on individual heads, and interpreting their role in “function application.”
Appendix C Details of datasets
@@
In addition to the country-capital task, we also conduct experiments on a product-developer task. Table 4 shows prompt templates we used. Here are Product-Developer ($X$, $Y$) pairs we used: (iPhone, Apple), (Windows7, Microsoft), (GTX1060, Nvidia), (YouTube, Google), (Firefox, Mozilla), (VirtualBox, Oracle), (Instagram, Facebook), (Pentium, Intel), (Steam, Valve), (Radeon, AMD), (Photoshop, Adobe), (PlayStation, Sony), (Kindle, Amazon), (GameBoy, Nintendo).
@@ 7 Keep in mind, the capital of $X$ is
8 Just a reminder, the capital of $X$ is
9 As we all know, the capital of $X$ is
10 According to the textbook, the capital of $X$ is
@@ 14 To emphasize, $X$ is developed by
Appendix D The influence of prompt templates
In experiments in the main text, the task word “capital” always precedes the country name $X$, as shown in Table 4. We now investigate whether altering the order of these keywords affects the factual recall. To explore this, we construct a new set of prompts, detailed in Table 5. We have the following key findings:
Table 5: Prompt templates in the country-capital task where $X$ precedes the task word “capital.”
Index Prompt Template
0 It’s crucial to know that $X$’s capital is
1 You are right to say that $X$’s capital is
2 Therefore, it’s correct to state that $X$’s capital is
3 When asked, always remember that $X$’s capital is
4 We confirm that $X$’s capital is
5 Don’t forget, $X$’s capital is
6 Bear in mind, $X$’s capital is
7 Keep in mind, $X$’s capital is
8 Just a reminder, $X$’s capital is
9 As we all know, $X$’s capital is
10 According to the textbook, $X$’s capital is
11 I am sure that $X$’s capital is
12 Without a doubt, $X$’s capital is
13 In case you didn’t know, $X$’s capital is
14 To emphasize, $X$’s capital is
Figure 10: Probability dynamics and detected influential heads across varied prompt templates.
(1) Consistent influential heads: Comparing Figure 10(a) and Figure 10(b), we can see that despite template differences, crucial task-specific heads, such as L9H8 and L10H0, maintain significance. L10H7 and L11H10 continue to suppress the correct prediction. Notably, some heads become more influential when using new templates such as L11H2 and L7H8. L11H2 is a mover head, and L7H8 attends to “capital,” appearing to impact the formation of task semantics, which falls out of the scope of this paper.
(2) Consistent probability dynamics’ patterns: The probability dynamics pattern remains akin to the original when using the new templates. Specifically, at layer 9, $X$ and $Y$ become evident in the residual stream, while at layer 10, the probability of $Y$ surpasses that of $X$ for the first time. The cosine similarity between current $\textbf{b}^{10,\text{proj}}$ and ($\textbf{W}_{U}[Y]-\textbf{W}_{U}[X]$) is 0.944. Recall that this similarity stands at 0.946 when employing the original prompt templates. The consistency across various templates underscores our previous findings regarding MLP10’s function in utilizing the intercept to steer the direction of the residual stream.
Moreover, at layer 11, the model exhibits the anti-overconfidence mechanism, with the cosine similarity between current $\textbf{b}^{11}$ and $\mathbb{E}[\textbf{W}_{U}]$ measuring at 0.89. This value is also notably close to 0.933 observed when employing the original templates. As shown in Figure 11(a), the anti-overconfidence mechanism can be suppressed by the two techniques we proposed.
(3) Prompt templates impact zero-shot prediction confidence: The primary distinction observed when utilizing two sets of prompts resides in the probability value. For prompts where $X$ precedes “capital,” the probabilities of both $X$ and $Y$ are lower. As depicted in Figure 10, this divergence originates primarily from layer 9. At layer 9, with original prompts, the probability of $X$ could reach 50%, whereas with the new prompts, it decreases to a mere 5%.
Upon scrutinizing the top decoded tokens given the new prompts, we found that “located” emerges as the predominant choice for output at layer 9 across almost all test samples. However, adding prefixes “$X_{\text{ICL}}$’s capital is $Y_{\text{ICL}}$” before our new prompts, to create few-shot test samples, eliminates the disparity of probability dynamics between given the new prompts and the original ones. The results are shown in Figure 11(b) and (c).
It seems that our new prompts, where we place “$X$” before “capital,” exacerbate the model’s uncertainty regarding the output domain, resulting in the model favoring the word “located” as a safe output. This uncertainty is effectively eliminated by in-context learning, aligning with the findings in Section 4 and Appendix A. Understanding how in-context learning boosts the model’s confidence remains a complex question, which we defer to future research endeavors as mentioned in the “Limitation” section.
Figure 11: (a) Our proposed techniques, including applying attention masks and subtracting the intercept from the output, can suppress the anti-overconfidence mechanism in the last layer. (b) One-shot probability dynamics given original/new prompts. (c) Two-shot probability dynamics given original/new prompts.
Appendix E Identifying the MLP layer for “function application”
In our main text, for simplicity’s sake, we followed Merullo et al. [21] to identify the layer where the probability of $Y$ surpasses that of $X$ as the point of function application. However, this principle is not always accurate. A more general principle would involve examining the layers “around” the point where the probabilities of $X$ and $Y$ begin to diverge significantly. To illustrate this, Figure 12 showcases the zero-shot probabilities of $X$ and $Y$ across the layers in GPT-2 medium and GPT-2 large models.
Appendix F GPT-2 small on the “product-developer” task
Figure 13: We test GPT-2 small on the product-developer task. (a) The probability of $X$ and $Y$ when early unembed each layer’s outputs. (b) The effect of patching path $\textbf{a}^{l,h}_{t=-1}\rightarrow\textbf{r}^{11}_{\text{post}}$. (c) Value weighted attention pattern of L10H0. (d) L10H0 is a mover head.
Appendix G Experiments on OPT-1.3B
OPT-1.3B [45] consists of 24 layers with 32 heads per layer. In Figure 16, we present the probability dynamics across layers, along with path patching experiments. From Figure 16(a), the argument passing is observed in layer 19. Figure 16(b) reveals that L19H15 in the zero-shot settings acts as the first influential argument passer. While L17H8 and L18H7 also serve as argument passers, their impact on increasing the probability of $X$ is diminished within the layer. Their activation is enhanced in few-shots scenarios, making the probability of $Y$ emerge in the residual stream early, as shown in Figure 19(b) and (c), which is similar to the behavior of L8H11 in GPT-2 small.
Recall that in Section 3.4, two key factors prompted us to realize that $\textbf{b}^{10}$ directs from $X$ to $Y$ within the MLP10 of GPT-2 small: Firstly, $X$ and $Y$ are equally important tokens that occupy most of the probability distribution. Secondly, after MLP10, the probability of $X$ decreases, but that of $Y$ increases. We have emphasized that these two factors do not always occur, and interpreting the mechanisms of MLPs requires a case-by-case analysis, depending on the model, task, or even prompts.
Table 6: Probability of several tokens decoded from certain nodes in OPT-1.3B.
Token Node $\textbf{r}^{20}_{\text{mid}}$ $\textbf{r}^{20}_{\text{post}}$
not 1.32% 4.97%
London 4.17% 3.21%
England 67.40% 74.40%
Regarding the final layer, as illustrated in Figure 19, OPT-1.3B also demonstrates anti-overconfidence tendencies, suppressing accurate predictions across zero-shot, one-shot, and two-shot scenarios. By employing the anti-suppression techniques outlined in Section 4, we effectively alleviate the anti-overconfidence mechanism employed by OPT.
Appendix H Experiments on Llama-2-7B and Llama-2-7B-chat
This section extends our experiments to another popular model family: Llama [36], specifically targeting Llama-2-7B and Llama-2-7B-chat. Llama-2-7B comprises 32 layers with 32 heads per layer. Llama-2-7B-chat is built upon Llama-2-7B, optimized on dialogue data through instruction fine-tuning [40] and RLHF (Reinforcement Learning from Human Feedback [48]). We present the experimental results of both models in Figure 17, confirming the validity of the mechanisms above. Moreover, we observed behavioral differences between the two models, offering insights into the impact of instruction tuning and RLHF. Figure 17 only presents the zero-shot results, as few-shot experiments yield similar conclusions.
Similar mechanisms shared in two models
The two models exhibit similar mechanisms for factual recall tasks: As shown in Figure 17(a) and (b), the “argument passing” process occurs at layer 19 in both models. From Figure 17(c) and (d), some mover heads are identified, such as L19H8 and L21H15. Two layers are crucial for “function application”: layer 19 and layer 22. In layer 19, the probability of $X$ soars for the first time but the MLP19 makes efforts to output $Y$. This is reflected in the corresponding cosine similarity values of 0.8822 and 0.8156 between $\textbf{b}^{19,\text{proj}}$ and $\textbf{W}_{U}[Y]-\textbf{W}_{U}[X]$ in two respective models. Meanwhile, in layer 22, where the cosine similarity values are 0.7722 and 0.8637 between $\textbf{b}^{22,\text{proj}}$ and $\textbf{W}_{U}[Y]-\textbf{W}_{U}[X]$, the decline of probability of $X$ and the increase of probability of $Y$ are interpreted. The MLP in the final layer exhibits the anti-overconfidence mechanism, which can be mitigated by subtracting the respective $\textbf{b}^{31,\text{proj}}$ from the residual stream, as shown in Figure 17(e) and (f). The behavior of the attention heads in the last layer will be discussed below.
Differences: base model versus chat model
Although both models demonstrate similar factual recall mechanisms, they exhibit divergent behaviors in certain aspects. Firstly, Llama-2-7B-chat displays notably higher confidence in achieving a probability of 76.60% to output $Y$ at the end, in contrast to Llama-2-7B, which yields a probability of only 29.11%. Secondly, attention heads in the final layer of Llama-2-7B (and -chat) positively contribute to accurate predictions, which contrasts with smaller models. However, MLPs in the last layer still suppress outputting $Y$ in a manner akin to other models.
One potential explanation for the first disparity is the impact of instruct tuning and RLHF on language models. These techniques aim to align the model’s output with human preferences, thereby potentially reducing the anti-overconfidence mechanism in chat models’ final layer (as it is more sensible for humans to respond given prompts with the capital city’ name rather than safe responses). Further investigation is necessary to fully comprehend the influence of instruct tuning and RLHF on language model mechanisms, as no prior research has explored this area to our knowledge. The second disparity stems from the nuanced interaction between active and inactive attention heads in the final layer, with our analysis revealing that only one attention head, L31H8, remains largely inactive, focusing its attention primarily on the initial token—a departure from behaviors observed in smaller models. Consequently, the collective effect of attention heads in the final layer exhibits positively.
Figure 17: Experimental results on Llama-2-7B and Llama-2-7B-chat. The argument-passing mechanism and the anti-overconfidence phenomenon are observed. Meanwhile, certain behaviors in the chat model differ slightly from those in all other base models studied in this paper.
Appendix I Details on solving the linear regression